OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? x+3y=5 4x+ ? y= ?

OpenStudy (anonymous):

They give you x+3y = 5, and the second equation is 4x, So multiply the first equation by 4

OpenStudy (anonymous):

You should get 4x+12y=20

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Can u help with one more plz?

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

What are the classifications of each system? Column A ---------------- 1. x+5y=-2 x+5y=4 2. y=3x+4 -2x+y=4 3. 5x+2y = -4 5x+2y=3 4. y= -x+3 2x+2y=6 --------------- Column B ---------------- A. consistent independent B. coincident C. inconsistent

OpenStudy (anonymous):

I'm not too sure. Sorry.

OpenStudy (anonymous):

Its ok thx tho